The probability that a vehicle entering the Luray Caverns has Canadian 

license plates is 0.12; the probability that it is a camper is 0.28; and the 

probability that it is a camper with Canadian license plates is 0.09. What is 

the probability that 

(i) A camper entering the Luray Caverns has Canadian license Plates.

(ii) A vehicle with Canadian license plates entering the Luray Caverns is a 

camper. 

(iii) A vehicle entering the Luray Caverns does not have Canadian plates or 

is not a camper.


1
Expert's answer
2021-06-15T08:54:11-0400

Solution

Let C denote the probability that a vehicle entering the Cavern has Canadian licence plates

"P(C)=0.12"

Let R denote the probability that it is a camper

"P(R) = 0.28"

The probability that it is a camper with Canadian plate license is

"P(C\bigcap R)= 0.09"


I) probability that a camper entering the Lurray has Canadian license plates

"P(C/R) = \frac{P(C\bigcap R)}{P(R)}"

"\frac{0.09}{0.28} = 0.321"


ii) To find probability that a vehicle with Canadian License plates entering the Lurray Cavern is a camper

we need to find;

"P(R/C) = \frac{P(R\bigcap C)}{P(C)}"

"\frac{0.09}{0.12} = 0.75"


iii) Probability that a vehicle entering the Luray Cavern does not have Canadian plates and is not a camper, We need to find

"P(C'\bigcup R')"

"1-P(C\bigcap R)"

"1-0.09"

"=0.91"